# Meta-Learning an In-Context Transformer Model of Human Higher Visual Cortex

### v1

@@ 4.2 Effectiveness of In-Context Higher Visual Cortex Encoders
On NSD. In this experiment, we evaluated BraInCoRL with each subject’s 9,000 unique images as in-context support set and the 1,000 shared images as test set, reporting explained variance averaged over category-selective ROIs in Table 1. We compare BraInCoRL using just 100 test images against within-subject ridge regression baseline trained on 100 and 300 support-set images of the test-subject, with the regularization strength selected via 5-fold cross-validation over $[10^{-3},10^{-2},\dots,10^{8}]$. Remarkably, BraInCoRL with only 100 images nearly matches the performance of the fully supervised reference model that is trained by gradient descent on each subject’s entire within-subject support-set of 9,000 images until convergence. We also evaluate an anatomical FsAverage baseline which aligns each training subject’s anatomy to a common template and projects the average response onto novel subjects for prediction. While this baseline benefits from a strong anatomical prior, it is outperformed by BraInCoRL, which directly adapts to each subject’s unique neural responses with higher efficiency.
Figure 4: UMAP visualization of predicted response weights. We apply UMAP to BraInCoRL -predicted voxelwise weights (100 support images) and show: (a) a flatmap for S1 with ROI outlines, (b) the same projection on an inflated surface, and (c) flatmaps for S2, S5, and S7. Color-coded clusters align with body/face regions (EBA, FFA/aTL-faces), place regions (RSC, OPA, PPA), and food regions (in red).
@@ 4.3 Semantic Discovery through Text–Image Embedding Alignment
Figure 6: Top contributing support images for each category-selective region in S1. For each of the five category-selective regions, we select the in-context support images with the highest attention weights in BraInCoRL’s final attention layer for voxels in that region. We visualize the top 5 contributing images for the place, word, face and body regions, and the top 10 for the food region.
Figure 7: Impact of support-set specificity on category-selective ROI encoding performance on NSD. We construct in-context support sets of 100 images based on descending semantic relevance for each ROI (tiers: 1–100, 101–200, 201–300, 301–400) and compare them with randomly sampled sets of equal size. Mean explained variance in the target category-selective ROIs increases as semantic specificity decreases, with all curated sets performing worse than random sampling. This suggests that overly specific support sets hinder generalization in voxelwise encoding. This pattern echoes prior findings on diverse stimuli contributing to better encoders [13].
@@ 5 Discussion
Here we have shown that meta-learning an in-context model of higher visual cortex can yield high performance and strong data efficiency gains, outperforming anatomical (FsAverage) and subject-wise baselines on novel subjects. Our work currently focuses on static natural images, and extensions to dynamic stimuli would likely require a rethinking of the encoder backbone and network structure. Further, while we show strong generalization across scanners and utilize the largest fMRI dataset that is NSD, there may still be limitations in dataset diversity [115]. Collection of larger and more diverse fMRI datasets will help mitigate this issue.
@@ Appendix A Technical Appendices and Supplementary Material
  3. 3.
@@
A.1 Implementation details
@@
    Input projection. We concatenate each image embedding with its scalar neural response and pass it through a single-layer MLP to align the two modalities into a unified internal feature space.
@@
    Transformer encoder. A stack of 20 self-attention layers integrates information across all support examples (plus learnable tokens), capturing contextual relationships and the relative importance of each in-context sample. We adopt best practices and utilize SwiGLU activation paired with pre-normalization in the attention block. We utilize 10 heads in each multi-head self attention.
@@
    Weight prediction. The aggregated representation is fed through another single-layer MLP that outputs a weight vector, which is then used to linearly combine unknown-image embeddings to produce the final neural response predictions.
The CLIP-based variant (encoding dimension $E=512$) contains approximately 97.2 M parameters; DINO ($E=768$) and SIGLIP ($E=1152$) variants comprise roughly 112 M and 130 M parameters, respectively.
@@
A.2 Text prompts for category-selective regions
@@
A.3 Cortex prediction explained variance for different image encoding backbones and subjects
@@
A.4 Voxelwise performance across five category-selective regions for different image encoding backbones and subjects
In this section, we report voxel-wise explained variance in five category-selective regions (faces, places, bodies, words, and food) with CLIP, DINO and SigLIP backbone for subjects S1-S8. We compare our in-context model (BraInCoRL) against the fully trained reference model fit to converge on each subject’s full 9,000-image training set, within-subject ridge regression baselines trained on 100 and 300 images, and the FsAverage map. BraInCoRL outperforms the ridge baselines and closely approaches the performance of the fully trained model.
Table S.1: Voxel-wise explained variance with the CLIP backbone for Subjects 1 and 2. We report performance for our in-context model (BraInCoRL), the fully trained reference (“Fully Trained”), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category-selective regions (faces, places, bodies, words, food).
@@
Table S.2: Voxel-wise explained variance with the CLIP backbone for Subjects 3 and 4. We report performance for our in-context model (BraInCoRL), the fully trained reference (“Fully Trained”), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category-selective regions (faces, places, bodies, words, food).
@@
Table S.3: Voxel-wise explained variance with the CLIP backbone for Subjects 5 and 6. We report performance for our in-context model (BraInCoRL), the fully trained reference ("Fully Trained"), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category-selective regions (faces, places, bodies, words, food).
@@
Table S.4: Voxel-wise explained variance with the CLIP backbone for Subjects 7 and 8. We report performance for our in-context model (BraInCoRL), the fully trained oracle ("Fully Trained"), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category-selective regions (faces, places, bodies, words, food).
@@
Table S.5: Voxel-wise explained variance with the DINO backbone for Subjects 1 and 2. We report performance for our in-context model (BraInCoRL), the fully trained oracle (“Fully Trained”), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category-selective regions (faces, places, bodies, words, food).
@@
Table S.6: Voxel-wise explained variance with the DINO backbone for Subjects 3 and 4. We report performance for our in-context model (BraInCoRL), the fully trained oracle (“Fully Trained”), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category-selective regions (faces, places, bodies, words, food).
@@
Table S.7: Voxel-wise explained variance with the DINO backbone for Subjects 5 and 6. We report performance for our in-context model (BraInCoRL), the fully trained reference ("Fully Trained"), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category-selective regions (faces, places, bodies, words, food).
@@
Table S.8: Voxel-wise explained variance with the DINO backbone for Subjects 7 and 8. We report performance for our in-context model (BraInCoRL), the fully trained oracle ("Fully Trained"), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category-selective regions (faces, places, bodies, words, food).
@@
Table S.9: Voxel-wise explained variance with the SigLIP backbone for Subjects 1 and 2. We report performance for our in-context model (BraInCoRL), the fully trained oracle (“Fully Trained”), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category-selective regions (faces, places, bodies, words, food).
@@
Table S.10: Voxel-wise explained variance with the SigLIP backbone for Subjects 3 and 4. We report performance for our in-context model (BraInCoRL), the fully trained oracle (“Fully Trained”), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category-selective regions (faces, places, bodies, words, food).
@@
Table S.11: Voxel-wise explained variance with the SigLIP backbone for Subjects 5 and 6. We report performance for our in-context model (BraInCoRL), the fully trained reference ("Fully Trained"), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category-selective regions (faces, places, bodies, words, food).
@@
Table S.12: Voxel-wise explained variance with the SigLIP backbone for Subjects 7 and 8. We report performance for our in-context model (BraInCoRL), the fully trained oracle ("Fully Trained"), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category-selective regions (faces, places, bodies, words, food).
@@
A.5 Voxelwise explained variance across varying support set sizes for more subjects, backbones and for pretrain-only models
In this section, we investigate how the size of the in-context support set affects voxelwise predictive performance. We evaluate three image-encoding backbones (CLIP, DINO, SigLIP) on eight subjects (S1–S8) by comparing the performance of BraInCoRL with the pretrain-only BraInCoRL (i.e. same architecture but only pretrained), the within-subject ridge-regression baseline, and the fully trained reference model fit to converge on each subject’s full 9,000-image training set.
@@
A.6 Impact of holding out the test subject’s unique images during meta-training evaluated on more backbones
@@
A.7 Correlation of each backbone’s predictions with fully trained activation predictions
Using various image-encoding backbones, we plot how each subject’s BraInCoRL predicted explained variance correlates with the fully trained model’s explained variance (fully trained model refers to the fully trained reference model fit to converge on each subject’s full 9,000-image training set). Across every backbone and all subjects, this correlation remains uniformly high.
Figure S.11: Voxelwise explained-variance correlation across backbones for subject 1, 2, 3, 4. Each panel shows a scatter of the BraInCoRL’s explained-variance predictions (100 in-context examples) versus the fully trained reference model’s explained variance for each voxel. Rows correspond to subjects (1, 2, 3, 4) and columns to image-encoding backbones (CLIP, DINO, SigLIP). The dashed line marks $y=x$.
Figure S.12: Voxelwise explained-variance correlation across backbones for subject 5, 6, 7, 8. Each panel shows a scatter of the BraInCoRL’s explained-variance predictions (100 in-context examples) versus the fully trained reference model’s explained variance for each voxel. Rows correspond to subjects (5, 6, 7, 8) and columns to image-encoding backbones (CLIP, DINO, SigLIP). The dashed line marks $y=x$.
A.8 Voxelwise explained-variance evaluation in BOLD5000 for more subjects and different backbones

### v2

@@ 4.2 Effectiveness of In-Context Higher Visual Cortex Encoders
On NSD. In this experiment, we evaluated BraInCoRL using each subject’s 9,000 unique images as the in-context support set and the shared 1,000 images as the test set. For each evaluation, we randomly sampled training images from the subject-specific in-context support set and test images from the shared test set. Explained variance statistics averaged over category-selective ROIs are reported in Table 1. We compare BraInCoRL using just 100 test images against within-subject ridge regression baseline trained on 100 and 300 support-set images of the test-subject, with the regularization strength selected via 5-fold cross-validation over $10^{-3},10^{-2},\dots,10^{8}]$. Remarkably, BraInCoRL with only 100 images nearly matches the performance of the fully supervised reference model that is trained by gradient descent on each subject’s entire within‐subject support-set of 9,000 images until convergence. We also evaluate an anatomical FsAverage baseline which aligns each training subject’s anatomy to a common template and projects the average response onto novel subjects for prediction. While this baseline benefits from a strong anatomical prior, it is outperformed by BraInCoRL, which directly adapts to each subject’s unique neural responses with higher efficiency.
Figure 4: UMAP visualization of predicted response weights. We apply UMAP to BraInCoRL -predicted voxelwise weights (100 support images) and show: (a) a flatmap for S1 with ROI outlines, (b) the same projection on an inflated surface, and (c) flatmaps for S2, S5, and S7. Color‐coded clusters align with body/face regions (EBA, FFA/aTL-faces), place regions (RSC, OPA, PPA), and food regions (in red).
@@ 4.3 Semantic Discovery through Text–Image Embedding Alignment
Figure 6: Top contributing support images for each category‐selective region in S1. For each of the five category‐selective regions, we select the in‐context support images with the highest attention weights in BraInCoRL’s final attention layer for voxels in that region. We visualize the top 5 contributing images for the place, word, face and body regions, and the top 10 for the food region.
Figure 7: Impact of support-set specificity on category-selective ROI encoding performance on NSD. We construct in-context support sets of 100 images based on descending semantic relevance for each ROI (tiers: 1–100, 101–200, 201–300, 301–400) and compare them with randomly sampled sets of equal size. Mean explained variance in the target category-selective ROIs increases as semantic specificity decreases, with all curated sets performing worse than random sampling. This suggests that overly specific support sets hinder generalization in voxelwise encoding. This pattern echoes prior findings on diverse stimuli contributing to better encoders [wang2022incorporating].
@@ 5 Discussion
Here we have shown that meta-learning an in-context model of higher visual cortex can yield high performance and strong data efficiency gains, outperforming anatomical (FsAverage) and subject-wise baselines on novel subjects. Our work currently focuses on static natural images, and extensions to dynamic stimuli would likely require a rethinking of the encoder backbone and network structure. Further, while we show strong generalization across scanners and utilize the largest fMRI dataset that is NSD, there may still be limitations in dataset diversity [shirakawa2024spurious]. Collection of larger and more diverse fMRI datasets will help mitigate this issue.
@@ Appendix A Technical Appendices and Supplementary Material
  4. 4.
    A more detailed description of logit scaling (Section A.4)
  5. 5.
@@
  14. 14.
    Additional evaluation of BraInCoRL on NSD dataset (Section A.14)
  15. 15.
    Additional evaluation of BraInCoRL on BOLD5000 dataset (Section A.15)
  16. 16.
    Evaluation of each training stage’s contribution (Section A.16)
  17. 17.
    Performance of BraInCoRL conditioned on the full 9000-image set (Section A.17)
  18. 18.
    Evaluation on the choice of loss function during training (Section A.18)
  19. 19.
    Ablation on logit scaling (Section A.19)
A.1 Social impacts
Our work introduces BraInCoRL, a meta‐learning framework that uses fMRI‐measured voxel responses and trained visual encoders to perform in‐context adaptation: given a small support set of image–response pairs, the model directly estimates each voxel’s response‐function parameters for novel stimuli. BraInCoRL’s fusion of diverse image data and neural measurements uncovers data‐driven principles of cortical organization beyond both traditional hypothesis‐driven experiments and computational encoding models that require thousands of samples. Moreover, its alignment of neural responses with natural language prompts enables the generation of new hypotheses about semantic representation in visual cortex. As such, BraInCoRL may accelerate early diagnosis and monitoring of visual or neurological disorders via rapid, subject‐specific cortical mapping; guide more efficient experimental design through optimized stimulus selection; deepen our understanding of semantic coding and inter‐subject variability; and, when integrated with generative image models, open avenues for brain‐guided stimulus synthesis and personalized neuroprosthetic and brain–computer interface development. While BraInCoRL offers significant potential for neuroscience and clinical applications, its reliance on fMRI datasets and computational infrastructure may limit accessibility for under-resourced research groups and raise privacy concerns if applied to sensitive neural data.
A.2 Implementation details
@@
    Input projection. An input projection MLP layer is applied to each token individually, which maps the stimulus semantics and voxel activation into an embedding. In detail, we concatenate each image embedding with its scalar neural response and pass it through a single‐layer MLP to align the two modalities into a unified internal feature space.
@@
    Transformer encoder. A stack of 20 self‐attention layers integrates information across all support examples (plus learnable tokens), capturing contextual relationships and the relative importance of each in‐context sample. We adopt best practices and utilize SwiGLU activation paired with pre-normalization in the attention block. We utilize 10 heads in each multi-head self attention.
@@
    Weight prediction. The [CLS] token from the final layer goes through an MLP to output a hyperweight which is used to parameterize the final encoder. In detail, the aggregated representation is fed through another single‐layer MLP that outputs a weight vector, which is then used to linearly combine unknown‐image embeddings to produce the final neural response predictions.
The CLIP‐based variant (encoding dimension $E=512$) contains approximately 97.2 M parameters; DINO ($E=768$) and SIGLIP ($E=1152$) variants comprise roughly 112 M and 130 M parameters, respectively.
@@
A.3 Text prompts for category-selective regions
@@
A.4 A more detailed description of logit scaling
The motivating factor underlying logit scaling is our desire to have our in-context learned encoder perform well regardless of the number of stimuli given to the model, and effectively generalize to context sizes beyond those seen during training. For example, while we may only train with between $30$ and $500$ images, a third-party experimenter may want to use fewer than $30$ images or more than $500$ images to condition the model. Across all cases, we want the model to succeed.
This logit scaling method was first proposed in [kexuefm-8823, chiang2022overcoming], and later adapted in Qwen LLM (logn-scaling) [bai2023qwen] and Llama 4 LLM (temperature scaling) [meta2025llama]. We will briefly summarize the high-level math, which we take from [kexuefm-8823] with our commentary:
Let the attention value in self-attention for a particular query token $i$ to value token $j$ to be
$a_{i,j}=\frac{e^{\lambda\mathbf{q}_{i}\cdot\mathbf{k}_{j}}}{\sum_{j=1}^{n}e^{\lambda\mathbf{q}_{i}\cdot\mathbf{k}_{j}}}$
Then the entropy is defined as
$\mathcal{H}_{i}=-\sum_{j=1}^{n}a_{i,j}\log a_{i,j}$
Substituting the expression for $a_{i,j}$ we have the entropy as
$\mathcal{H}_{i}=-\sum_{j=1}^{n}a_{i,j}\log\left(\frac{e^{\lambda\mathbf{q}_{i}\cdot\mathbf{k}_{j}}}{\sum_{j=1}^{n}e^{\lambda\mathbf{q}_{i}\cdot\mathbf{k}_{j}}}\right)$
Since $\sum_{j=1}^{n}a_{i,j}=1$,
we can express the formula as
$\mathcal{H}_{i}=\log\sum_{j=1}^{n}e^{\lambda\mathbf{q}_{i}\cdot\mathbf{k}_{j}}-\frac{\sum\limits_{j=1}^{n}e^{\lambda\mathbf{q}_{i}\cdot\mathbf{k}_{j}}(\lambda\mathbf{q}_{i}\cdot\mathbf{k}_{j})}{\sum\limits_{j=1}^{n}e^{\lambda\mathbf{q}_{i}\cdot\mathbf{k}_{j}}}$
If the first term is expressed as an expectation over $j$, we have
$\sum_{j=1}^{n}e^{\lambda\mathbf{q}_{i}\cdot\mathbf{k}_{j}}=n\times\frac{1}{n}\sum_{j=1}^{n}e^{\lambda\mathbf{q}_{i}\cdot\mathbf{k}_{j}}\approx n\mathbb{E}_{j}[e^{\lambda\mathbf{q}_{i}\cdot\mathbf{k}_{j}}]$
Which leads to the following approximation of entropy
$\mathcal{H}_{i}\approx\log n+\log\mathbb{E}_{j}[e^{\lambda\mathbf{q}_{i}\cdot\mathbf{k}_{j}}]-\frac{\lambda\mathbb{E}_{j}[e^{\lambda\mathbf{q}_{i}\cdot\mathbf{k}_{j}}(\mathbf{q}_{i}\cdot\mathbf{k}_{j})]}{\mathbb{E}_{j}[e^{\lambda\mathbf{q}_{i}\cdot\mathbf{k}_{j}}]}$
If the vectors are assumed to be the output of a layernorm layer with length $\sqrt{d}$, the expectation can be converted to one over the angles between vectors
$\mathcal{H}_{i}\approx\log n+\log\mathbb{E}_{\theta}[e^{\lambda d\cos\theta}]-\frac{\lambda d\mathbb{E}_{\theta}[e^{\lambda d\cos\theta}\cos\theta]}{\mathbb{E}_{\theta}[e^{\lambda d\cos\theta}]}$
Since most randomly distributed vectors in higher dimensions are orthogonal, we derive a term which can roughly be expressed as
$\mathcal{H}_{i}\approx\log n+C$
where $C$ does not depend on the number of tokens $n$.
This leads to an approximate scaling factor for the logits of $\log n$ to keep the entropy invariant to context length.
Therefore, we change the standard formulation of attention values by applying a scaling factor of $\log n$ to the $QK^{T}$ term, as shown in Equation (4) of our paper.
Note that in the above explanation we adopt the notation from [kexuefm-8823].
A.5 Cortex prediction explained variance for different image encoding backbones and subjects
@@
A.6 Voxelwise performance across five category-selective regions for different image encoding backbones and subjects
In this section, we report voxel‐wise explained variance in five category‐selective regions (faces, places, bodies, words, and food) with CLIP, DINO and SigLIP backbone for subjects S1-S8. We compare our in‐context model (BraInCoRL) against the fully trained reference model fit to converge on each subject’s full 9,000-image training set, within-subject ridge regression baselines trained on 100 and 300 images, and the FsAverage map. BraInCoRL outperforms the ridge baselines and closely approaches the performance of the fully trained model.
Table S.1: Voxel‐wise explained variance with the CLIP backbone for Subjects 1 and 2. We report performance for our in‐context model (BraInCoRL), the fully trained reference (“Fully Trained”), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category‐selective regions (faces, places, bodies, words, food).
@@
Table S.2: Voxel‐wise explained variance with the CLIP backbone for Subjects 3 and 4. We report performance for our in‐context model (BraInCoRL), the fully trained reference (“Fully Trained”), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category‐selective regions (faces, places, bodies, words, food).
@@
Table S.3: Voxel‐wise explained variance with the CLIP backbone for Subjects 5 and 6. We report performance for our in‐context model (BraInCoRL), the fully trained reference (”Fully Trained”), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category‐selective regions (faces, places, bodies, words, food).
@@
Table S.4: Voxel‐wise explained variance with the CLIP backbone for Subjects 7 and 8. We report performance for our in‐context model (BraInCoRL), the fully trained oracle (”Fully Trained”), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category‐selective regions (faces, places, bodies, words, food).
@@
Table S.5: Voxel‐wise explained variance with the DINO backbone for Subjects 1 and 2. We report performance for our in‐context model (BraInCoRL), the fully trained oracle (“Fully Trained”), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category‐selective regions (faces, places, bodies, words, food).
@@
Table S.6: Voxel‐wise explained variance with the DINO backbone for Subjects 3 and 4. We report performance for our in‐context model (BraInCoRL), the fully trained oracle (“Fully Trained”), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category‐selective regions (faces, places, bodies, words, food).
@@
Table S.7: Voxel‐wise explained variance with the DINO backbone for Subjects 5 and 6. We report performance for our in‐context model (BraInCoRL), the fully trained reference (”Fully Trained”), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category‐selective regions (faces, places, bodies, words, food).
@@
Table S.8: Voxel‐wise explained variance with the DINO backbone for Subjects 7 and 8. We report performance for our in‐context model (BraInCoRL), the fully trained oracle (”Fully Trained”), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category‐selective regions (faces, places, bodies, words, food).
@@
Table S.9: Voxel‐wise explained variance with the SigLIP backbone for Subjects 1 and 2. We report performance for our in‐context model (BraInCoRL), the fully trained oracle (“Fully Trained”), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category‐selective regions (faces, places, bodies, words, food).
@@
Table S.10: Voxel‐wise explained variance with the SigLIP backbone for Subjects 3 and 4. We report performance for our in‐context model (BraInCoRL), the fully trained oracle (“Fully Trained”), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category‐selective regions (faces, places, bodies, words, food).
@@
Table S.11: Voxel‐wise explained variance with the SigLIP backbone for Subjects 5 and 6. We report performance for our in‐context model (BraInCoRL), the fully trained reference (”Fully Trained”), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category‐selective regions (faces, places, bodies, words, food).
@@
Table S.12: Voxel‐wise explained variance with the SigLIP backbone for Subjects 7 and 8. We report performance for our in‐context model (BraInCoRL), the fully trained oracle (”Fully Trained”), within-subject ridge regression baselines (100, 300), and the FsAverage map across five category‐selective regions (faces, places, bodies, words, food).
@@
A.7 Voxelwise explained variance across varying support set sizes for more subjects, backbones and for pretrain-only models
In this section, we investigate how the size of the in‐context support set affects voxelwise predictive performance. We evaluate three image‐encoding backbones (CLIP, DINO, SigLIP) on eight subjects (S1–S8) by comparing the performance of BraInCoRL with the pretrain-only BraInCoRL (i.e. same architecture but only pretrained), the within-subject ridge-regression baseline, and the fully trained reference model fit to converge on each subject’s full 9,000-image training set.
@@
A.8 Impact of holding out the test subject’s unique images during meta-training evaluated on more backbones
@@
A.9 Correlation of each backbone’s predictions with fully trained activation predictions
Using various image‐encoding backbones, we plot how each subject’s BraInCoRL predicted explained variance correlates with the fully trained model’s explained variance (fully trained model refers to the fully trained reference model fit to converge on each subject’s full 9,000-image training set). Across every backbone and all subjects, this correlation remains uniformly high.
Figure S.11: Voxelwise explained‐variance correlation across backbones for subject 1, 2, 3, 4. Each panel shows a scatter of the BraInCoRL’s explained‐variance predictions (100 in‐context examples) versus the fully trained reference model’s explained variance for each voxel. Rows correspond to subjects (1, 2, 3, 4) and columns to image‐encoding backbones (CLIP, DINO, SigLIP). The dashed line marks $y=x$.
Figure S.12: Voxelwise explained‐variance correlation across backbones for subject 5, 6, 7, 8. Each panel shows a scatter of the BraInCoRL’s explained‐variance predictions (100 in‐context examples) versus the fully trained reference model’s explained variance for each voxel. Rows correspond to subjects (5, 6, 7, 8) and columns to image‐encoding backbones (CLIP, DINO, SigLIP). The dashed line marks $y=x$.
A.10 Voxelwise explained-variance evaluation in BOLD5000 for more subjects and different backbones